OpenStudy (zned6559):

A certain reaction produces 20.7 L of gas at 1.00 atm of constant pressure. If 1350 J of heat is absorbed by the reaction what is the change in internal energy (ΔU)?

OpenStudy (zned6559):

I know that delta U = q+w and w=-p delta v

OpenStudy (aaronq):

It's straight forward if you can assign signs (negative/positive) to the values. is work done or by the system? is heat absorbed or released?

OpenStudy (zned6559):

delta u = 1350 J - 20.7 = 1329.3, but this is the wrong answer

OpenStudy (zned6559):

the correct answer is -747J

OpenStudy (aaronq):

it has to do with the units, you gotta convert the pressure, L*atm is not equal to J

OpenStudy (aaronq):

divide the work by (9.8692*10^-3)

OpenStudy (zned6559):

oh i got it now!

OpenStudy (aaronq):

awesome